In an ionic solution, 5.0×1015 positive ions with charge +2e pass to the right each second while 6.0×1015 negative ions with charge −e pass to the left.

What is the magnitude of current in the solution? (mA)

What is the direction of current in the solution? (left or right?)
Answer #1

n1 = no of positive ions = 5*1015

q1 = charge on positive ions = 2e

n2 = no of negative ions = 6*1015

q2 = charge on negative ions = - e

also e = 1.6*10-19 C

time, t = 1 sec

current , I = q / t

current will be in the direction of positive ions ie towards right.

so I = ( q1 + q2 ) / t

     I = ( 5* 1015 *2*1.6*10-19 + 6*1015*( - 1.6*10-19)*( -1) )                                                 ( q = ne)

[ multiply by (-1) because negative ions in opposite direction to that of positive ions)

I = 2.56 mA